### v3

@@ Appendix C Additional results
The remainder of this section contains additional results of the empirical application from Section 7. Table C.1 lists the unipolar markers of the three Big Five scales used by Arias et al., (2020), namely extroversion, conscientiousness, and neuroticism. Tables C.2C.4 contain the (polychoric) correlation matrices of the items in each scale, estimated by maximum likelihood and our robust estimator, while Figures C.2C.4 visualize the absolute difference between the two estimators for each pairwise correlation.
Figure C.1: Mean squared error of all estimated parameters in the polychoric model, $\|\widehat{\bm{\theta}}_{N}-\bm{\theta}_{*}\|^{2}$, in the first simulation in Section 6.
@@
Table C.1: Unipolar markers of three Big Five personality traits (Goldberg,, 1992). Each trait is measured by six pairs of items, where each item is a single English adjective. Each item pair consists of a positive and negative item. We explain item identifiers by means of the following example. Item “C3_P” refers to the positive (P) item in the 3rd pair of the conscientiousness (C) scale, that is, adjective “conscientious”, whereas “N1_N” would refer to “angry”.
N1_P N1_N N2_P N2_N N3_P N3_N N4_P N4_N N5_P N5_N N6_P N6_N
N1_P 1.00 -0.37 0.71 -0.50 0.69 -0.49 0.27 -0.24 0.58 -0.47 0.42 -0.32
N1_N -0.37 1.00 -0.40 0.55 -0.39 0.47 -0.19 0.40 -0.39 0.60 -0.32 0.56
N2_P 0.71 -0.40 1.00 -0.55 0.75 -0.54 0.26 -0.26 0.55 -0.41 0.53 -0.47
N2_N -0.50 0.55 -0.55 1.00 -0.54 0.65 -0.24 0.42 -0.41 0.57 -0.31 0.52
N3_P 0.69 -0.39 0.75 -0.54 1.00 -0.53 0.29 -0.28 0.63 -0.44 0.52 -0.48
N3_N -0.49 0.47 -0.54 0.65 -0.53 1.00 -0.28 0.43 -0.44 0.58 -0.29 0.47
N4_P 0.27 -0.19 0.26 -0.24 0.29 -0.28 1.00 -0.61 0.26 -0.20 0.18 -0.20
N4_N -0.24 0.40 -0.26 0.42 -0.28 0.43 -0.61 1.00 -0.33 0.46 -0.22 0.44
N5_P 0.58 -0.39 0.55 -0.41 0.63 -0.44 0.26 -0.33 1.00 -0.69 0.53 -0.46
N5_N -0.47 0.60 -0.41 0.57 -0.44 0.58 -0.20 0.46 -0.69 1.00 -0.35 0.57
N6_P 0.42 -0.32 0.53 -0.31 0.52 -0.29 0.18 -0.22 0.53 -0.35 1.00 -0.58
N6_N -0.32 0.56 -0.47 0.52 -0.48 0.47 -0.20 0.44 -0.46 0.57 -0.58 1.00
(a) Maximum likelihood estimates
@@
(b) Robust estimates
Table C.2: Estimated correlation matrices of the items in the neuroticism scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=0.6$ (bottom panel). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
E1_P E1_N E2_P E2_N E3_P E3_N E4_P E4_N E5_P E5_N E6_P E6_N
E1_P 1.00 -0.77 0.50 -0.26 0.70 -0.50 0.56 -0.42 0.51 -0.40 0.51 -0.32
E1_N -0.77 1.00 -0.38 0.34 -0.59 0.61 -0.45 0.54 -0.47 0.50 -0.35 0.37
E2_P 0.50 -0.38 1.00 -0.65 0.43 -0.27 0.49 -0.28 0.47 -0.38 0.54 -0.39
E2_N -0.26 0.34 -0.65 1.00 -0.24 0.34 -0.30 0.40 -0.32 0.48 -0.38 0.49
E3_P 0.70 -0.59 0.43 -0.24 1.00 -0.59 0.44 -0.36 0.46 -0.40 0.41 -0.25
E3_N -0.50 0.61 -0.27 0.34 -0.59 1.00 -0.27 0.56 -0.35 0.45 -0.24 0.37
E4_P 0.56 -0.45 0.49 -0.30 0.44 -0.27 1.00 -0.41 0.64 -0.49 0.54 -0.34
E4_N -0.42 0.54 -0.28 0.40 -0.36 0.56 -0.41 1.00 -0.49 0.60 -0.27 0.40
E5_P 0.51 -0.47 0.47 -0.32 0.46 -0.35 0.64 -0.49 1.00 -0.71 0.39 -0.23
E5_N -0.40 0.50 -0.38 0.48 -0.40 0.45 -0.49 0.60 -0.71 1.00 -0.34 0.45
E6_P 0.51 -0.35 0.54 -0.38 0.41 -0.24 0.54 -0.27 0.39 -0.34 1.00 -0.68
E6_N -0.32 0.37 -0.39 0.49 -0.25 0.37 -0.34 0.40 -0.23 0.45 -0.68 1.00
(a) Maximum likelihood estimates
@@
(b) Robust estimates
Table C.3: Estimated correlation matrices of the items in the extroversion scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=0.6$ (bottom panel). The items are “extraverted” (E1_P), “introverted” (E1_N), “energetic” (E2_P), “unenergetic” (E2_N), “talkative” (E3_P), “silent” (E3_N), “bold” (E4_P), “timid” (E4_N), “assertive” (E5_P), “unassertive” (E5_N), “adventurous” (E6_P), and “unadventurous” (E6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
C1_P C1_N C2_P C2_N C3_P C3_N C4_P C4_N C5_P C5_N C6_P C6_N
C1_P 1.00 -0.76 0.56 -0.42 0.34 -0.35 0.37 -0.26 0.51 -0.40 0.43 -0.43
C1_N -0.76 1.00 -0.51 0.58 -0.24 0.55 -0.32 0.44 -0.43 0.61 -0.44 0.55
C2_P 0.56 -0.51 1.00 -0.69 0.42 -0.55 0.57 -0.42 0.51 -0.54 0.65 -0.55
C2_N -0.42 0.58 -0.69 1.00 -0.39 0.75 -0.48 0.67 -0.42 0.70 -0.53 0.63
C3_P 0.34 -0.24 0.42 -0.39 1.00 -0.32 0.39 -0.34 0.44 -0.33 0.38 -0.25
C3_N -0.35 0.55 -0.55 0.75 -0.32 1.00 -0.37 0.59 -0.38 0.71 -0.44 0.53
C4_P 0.37 -0.32 0.57 -0.48 0.39 -0.37 1.00 -0.51 0.36 -0.38 0.39 -0.31
C4_N -0.26 0.44 -0.42 0.67 -0.34 0.59 -0.51 1.00 -0.38 0.59 -0.31 0.43
C5_P 0.51 -0.43 0.51 -0.42 0.44 -0.38 0.36 -0.38 1.00 -0.43 0.54 -0.39
C5_N -0.40 0.61 -0.54 0.70 -0.33 0.71 -0.38 0.59 -0.43 1.00 -0.43 0.53
C6_P 0.43 -0.44 0.65 -0.53 0.38 -0.44 0.39 -0.31 0.54 -0.43 1.00 -0.61
C6_N -0.43 0.55 -0.55 0.63 -0.25 0.53 -0.31 0.43 -0.39 0.53 -0.61 1.00
(a) Maximum likelihood estimates
@@
(b) Robust estimates
@@
Table C.4: Estimated correlation matrices of the items in the conscientiousness scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=0.6$ (bottom panel). The items are “calm” (C1_P), “angry” (C1_N), “relaxed” (C2_P), “tense” (C2_N), “at ease” (C3_P), “nervous” (C3_N), “not envious” (C4_P), “envious” (C4_N), “stable” (C5_P), “unstable” (C5_N), “contented” (C6_P), and “discontented” (C6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Figure C.2: Difference between absolute estimates for the polychoric correlation coefficient of our robust estimator and the MLE for each item pair in the neuroticism scale, using the data of Arias et al., (2020). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Figure C.3: Difference between absolute estimates for the polychoric correlation coefficient of our robust estimator and the MLE for each item pair in the extroversion scale, using the data of Arias et al., (2020). The items are “extraverted” (E1_P), “introverted” (E1_N), “energetic” (E2_P), “unenergetic” (E2_N), “talkative” (E3_P), “silent” (E3_N), “bold” (E4_P), “timid” (E4_N), “assertive” (E5_P), “unassertive” (E5_N), “adventurous” (E6_P), and “unadventurous” (E6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Figure C.4: Difference between absolute estimates for the polychoric correlation coefficient of our robust estimator and the MLE for each item pair in the conscientiousness scale, using the data of Arias et al., (2020). The items are “calm” (C1_P), “angry” (C1_N), “relaxed” (C2_P), “tense” (C2_N), “at ease” (C3_P), “nervous” (C3_N), “not envious” (C4_P), “envious” (C4_N), “stable” (C5_P), “unstable” (C5_N), “contented” (C6_P), and “discontented” (C6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
$X$\$Y$ 1 2 3 4 5
1 9,814,457,556.73 16,010.33 10.82 0.14 $-0.35$
2 2,423.07 9.07 $-0.20$ $-0.10$ 0.42
3 4.48 $-0.35$ $-0.01$ $-0.20$ 76.14
4 $-0.12$ $-0.08$ $-0.39$ 10.66 222,527.08
5 $-0.11$ $-0.12$ 35.01 55,419.33 995,017,243,196.60
Table C.5: Pearson residual, $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})-1$, of each response $(x,y)$ for the “not envious” ($X$)–“envious” ($Y$) item pair in the measurements of Arias et al., (2020) of a neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed with tuning constant $c=0.6$.
@@
Figure C.5: Dot plot of cells for the neuroticism item adjective pair “envious” and “not envious” in the data of Arias et al., (2020), where each item has five Likert-type response options, anchored by “very inaccurate” (= 1) and “very accurate” (= 5). Each dot’s size is proportional to the relative empirical frequency of its associated contingency table cell, $\widehat{f}_{N}(x,y)$, whereas its color varies by the value of the cell’s PR, $\widehat{f}_{N}(x,y)/p_{xy}(\widehat{\bm{\theta}}_{N})-1$, at robust parameter estimate with tuning constant $c=0.6$. Note that some cells have an extremely large PR (substantially larger than ideal value 0; see also Table C.5), indicating a poor model fit for those cells. We therefore censored the color scale so that a fully saturated red is assigned to all cells with PR $\geq 3$, i.e., the empirical frequency $\widehat{f}_{N}(x,y)$ being at least four times as large as the model frequency $p_{xy}(\widehat{\bm{\theta}}_{N})$.

### v4

@@ Appendix F Additional results for the empirical application
This section contains additional results of the empirical application from Section 7. Table F.1 lists the unipolar markers of the three Big Five scales used by Arias et al., (2020), namely extroversion, conscientiousness, and neuroticism. Tables F.2F.4 contain the (polychoric) correlation matrices of the items in each scale, estimated by maximum likelihood and our robust estimator, while Figures F.1F.3 visualize the absolute difference between the two estimators for each pairwise correlation. Furthermore, Table F.5 contains the cellwise Pearson residuals for the item pair “not envious” and “envious” in the neuroticism scale. Figure F.4 provides a corresponding visualization.
@@
Table F.1: Unipolar markers of three Big Five personality traits (Goldberg,, 1992). Each trait is measured by six pairs of items, where each item is a single English adjective. Each item pair consists of a positive and negative item. We explain item identifiers by means of the following example. Item “C3_P” refers to the positive (P) item in the 3rd pair of the conscientiousness (C) scale, that is, adjective “conscientious”, whereas “N1_N” would refer to “angry”.
@@
(a) Robust estimates
N1_P N1_N N2_P N2_N N3_P N3_N N4_P N4_N N5_P N5_N N6_P N6_N
N1_P 1.00 -0.37 0.71 -0.50 0.69 -0.49 0.27 -0.24 0.58 -0.47 0.42 -0.32
N1_N -0.37 1.00 -0.40 0.55 -0.39 0.48 -0.19 0.40 -0.39 0.60 -0.32 0.57
N2_P 0.71 -0.40 1.00 -0.55 0.75 -0.54 0.26 -0.26 0.55 -0.41 0.53 -0.47
N2_N -0.50 0.55 -0.55 1.00 -0.54 0.65 -0.24 0.42 -0.41 0.57 -0.31 0.52
N3_P 0.69 -0.39 0.75 -0.54 1.00 -0.53 0.29 -0.28 0.63 -0.44 0.52 -0.48
N3_N -0.49 0.48 -0.54 0.65 -0.53 1.00 -0.29 0.43 -0.45 0.58 -0.29 0.47
N4_P 0.27 -0.19 0.26 -0.24 0.29 -0.29 1.00 -0.62 0.26 -0.20 0.18 -0.20
N4_N -0.24 0.40 -0.26 0.42 -0.28 0.43 -0.62 1.00 -0.33 0.46 -0.22 0.44
N5_P 0.58 -0.39 0.55 -0.41 0.63 -0.45 0.26 -0.33 1.00 -0.70 0.53 -0.46
N5_N -0.47 0.60 -0.41 0.57 -0.44 0.58 -0.20 0.46 -0.70 1.00 -0.35 0.57
N6_P 0.42 -0.32 0.53 -0.31 0.52 -0.29 0.18 -0.22 0.53 -0.35 1.00 -0.58
N6_N -0.32 0.57 -0.47 0.52 -0.48 0.47 -0.20 0.44 -0.46 0.57 -0.58 1.00
(b) Maximum likelihood estimates
Table F.2: Estimated correlation matrices of the items in the neuroticism scale from the data in Arias et al., (2020, Sample 1, $N=725$) using the robust estimator with $c=0.6$ (top) and the MLE (bottom). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
(a) Robust estimates
E1_P E1_N E2_P E2_N E3_P E3_N E4_P E4_N E5_P E5_N E6_P E6_N
E1_P 1.00 -0.78 0.50 -0.26 0.70 -0.50 0.56 -0.42 0.51 -0.40 0.52 -0.32
E1_N -0.78 1.00 -0.38 0.34 -0.59 0.61 -0.45 0.54 -0.47 0.50 -0.35 0.37
E2_P 0.50 -0.38 1.00 -0.65 0.43 -0.27 0.49 -0.28 0.47 -0.38 0.55 -0.39
E2_N -0.26 0.34 -0.65 1.00 -0.24 0.34 -0.30 0.40 -0.32 0.48 -0.38 0.50
E3_P 0.70 -0.59 0.43 -0.24 1.00 -0.59 0.44 -0.36 0.46 -0.40 0.41 -0.25
E3_N -0.50 0.61 -0.27 0.34 -0.59 1.00 -0.27 0.56 -0.35 0.45 -0.24 0.37
E4_P 0.56 -0.45 0.49 -0.30 0.44 -0.27 1.00 -0.41 0.64 -0.49 0.54 -0.34
E4_N -0.42 0.54 -0.28 0.40 -0.36 0.56 -0.41 1.00 -0.49 0.60 -0.27 0.40
E5_P 0.51 -0.47 0.47 -0.32 0.46 -0.35 0.64 -0.49 1.00 -0.71 0.39 -0.23
E5_N -0.40 0.50 -0.38 0.48 -0.40 0.45 -0.49 0.60 -0.71 1.00 -0.34 0.45
E6_P 0.52 -0.35 0.55 -0.38 0.41 -0.24 0.54 -0.27 0.39 -0.34 1.00 -0.68
E6_N -0.32 0.37 -0.39 0.50 -0.25 0.37 -0.34 0.40 -0.23 0.45 -0.68 1.00
(b) Maximum likelihood estimates
Table F.3: Estimated correlation matrices of the items in the extroversion scale from the data in Arias et al., (2020, Sample 1, $N=725$) using the robust estimator with $c=0.6$ (top) and the MLE (bottom). The items are “extraverted” (E1_P), “introverted” (E1_N), “energetic” (E2_P), “unenergetic” (E2_N), “talkative” (E3_P), “silent” (E3_N), “bold” (E4_P), “timid” (E4_N), “assertive” (E5_P), “unassertive” (E5_N), “adventurous” (E6_P), and “unadventurous” (E6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
(a) Robust estimates
C1_P C1_N C2_P C2_N C3_P C3_N C4_P C4_N C5_P C5_N C6_P C6_N
C1_P 1.00 -0.77 0.56 -0.43 0.34 -0.35 0.38 -0.26 0.51 -0.41 0.43 -0.43
C1_N -0.77 1.00 -0.51 0.59 -0.24 0.55 -0.32 0.44 -0.43 0.61 -0.44 0.55
C2_P 0.56 -0.51 1.00 -0.70 0.42 -0.56 0.57 -0.43 0.51 -0.54 0.65 -0.55
C2_N -0.43 0.59 -0.70 1.00 -0.40 0.75 -0.48 0.68 -0.43 0.71 -0.53 0.63
C3_P 0.34 -0.24 0.42 -0.40 1.00 -0.32 0.39 -0.34 0.44 -0.34 0.38 -0.25
C3_N -0.35 0.55 -0.56 0.75 -0.32 1.00 -0.37 0.60 -0.38 0.72 -0.45 0.54
C4_P 0.38 -0.32 0.57 -0.48 0.39 -0.37 1.00 -0.52 0.36 -0.39 0.39 -0.31
C4_N -0.26 0.44 -0.43 0.68 -0.34 0.60 -0.52 1.00 -0.38 0.59 -0.31 0.43
C5_P 0.51 -0.43 0.51 -0.43 0.44 -0.38 0.36 -0.38 1.00 -0.43 0.54 -0.39
C5_N -0.41 0.61 -0.54 0.71 -0.34 0.72 -0.39 0.59 -0.43 1.00 -0.43 0.53
C6_P 0.43 -0.44 0.65 -0.53 0.38 -0.45 0.39 -0.31 0.54 -0.43 1.00 -0.61
C6_N -0.43 0.55 -0.55 0.63 -0.25 0.54 -0.31 0.43 -0.39 0.53 -0.61 1.00
(b) Maximum likelihood estimates
@@
Table F.4: Estimated correlation matrices of the items in the conscientiousness scale from the data in Arias et al., (2020, Sample 1, $N=725$) using the robust estimator with $c=0.6$ (top) and the MLE (bottom). The items are “calm” (C1_P), “angry” (C1_N), “relaxed” (C2_P), “tense” (C2_N), “at ease” (C3_P), “nervous” (C3_N), “not envious” (C4_P), “envious” (C4_N), “stable” (C5_P), “unstable” (C5_N), “contented” (C6_P), and “discontented” (C6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Figure F.1: Difference between absolute estimates for the polychoric correlation coefficient of the robust estimator with $c=0.6$ and the MLE for each item pair in the neuroticism scale, using the data of Arias et al., (2020). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Figure F.2: Difference between absolute estimates for the polychoric correlation coefficient of the robust estimator with $c=0.6$ and the MLE for each item pair in the extroversion scale, using the data of Arias et al., (2020). The items are “extraverted” (E1_P), “introverted” (E1_N), “energetic” (E2_P), “unenergetic” (E2_N), “talkative” (E3_P), “silent” (E3_N), “bold” (E4_P), “timid” (E4_N), “assertive” (E5_P), “unassertive” (E5_N), “adventurous” (E6_P), and “unadventurous” (E6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Figure F.3: Difference between absolute estimates for the polychoric correlation coefficient of the robust estimator with $c=0.6$ and the MLE for each item pair in the conscientiousness scale, using the data of Arias et al., (2020). The items are “calm” (C1_P), “angry” (C1_N), “relaxed” (C2_P), “tense” (C2_N), “at ease” (C3_P), “nervous” (C3_N), “not envious” (C4_P), “envious” (C4_N), “stable” (C5_P), “unstable” (C5_N), “contented” (C6_P), and “discontented” (C6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
[…]
1 9,780,473,685.31 15,982.37 10.81 0.14 0.35-0.35- 0.35
2 2,419.15 9.06 0.20-0.20- 0.20 0.10-0.10- 0.10 0.42
3 4.48 0.35-0.35- 0.35 0.01-0.01- 0.01 0.20-0.20- 0.20 76.11
4 0.12-0.12- 0.12 0.08-0.08- 0.08 0.39-0.39- 0.39 11.66 222,240.11
5 0.11-0.11- 0.11 0.12-0.12- 0.12 34.98 55,329.21 991,790,294,422.36
[…]
@@
Figure F.4: Dot plot of cells for the neuroticism item adjective pair “not envious” and “envious” in the data of Arias et al., (2020), where each item has five Likert-type response options, anchored by “very inaccurate” (= 1) and “very accurate” (= 5). Each dot’s size is proportional to the relative empirical frequency of its associated contingency table cell, $\widehat{f}_{N}(x,y)$, whereas its color varies by the value of the cell’s PR, $\widehat{f}_{N}(x,y)/p_{xy}(\widehat{\bm{\theta}}_{N})-1$, at robust parameter estimate with tuning constant $c=0.6$. Note that some cells have an extremely large PR (substantially larger than the ideal value 0; see also Table F.5), indicating a poor model fit for those cells. We therefore censored the color scale so that a fully saturated red is assigned to all cells with PR 3$\geq$ 3≥ 3, i.e., the empirical frequency $\widehat{f}_{N}(x,y)$ being at least four times as large as the model frequency $p_{xy}(\widehat{\bm{\theta}}_{N})$.